OpenStudy (anonymous):

Help! I am in an astronomy course and we are learning the small angular formula...I am so lost, =/

OpenStudy (unklerhaukus):

for a small angle \(\theta\) \(h\sim a\) 

OpenStudy (anonymous):

 The small angle approximation will also be useful, where:\[\sin \theta = \theta\] so\[b= \theta h, or \theta a, due \to the \small angular formula\]